OpenStudy (anonymous):

Solve 2x2 + 12x − 14 = 0 by completing the square. 2(x + 6)2 = 26; x = −6 + square root 13, x = −6 − square root 13 2(x + 6)2 = 20; x = −6 + square root 20, x = −6 − square root 20 2(x + 3)2 = 32; x = 1, x = −7 2(x + 3)2 = 32; x = 13, x = 19

TheSmartOne (thesmartone):

So do you want someone to check your answers?

TheSmartOne (thesmartone):

\[2x^2+12x-14=0\] \[x^2+6x-7=0\] \[x^2+6x=7\] \[x^2+6x+9=7+9\] \[(x+3)^2=16\] \[x+3=\pm4\] \[x=-3\pm4\]

TheSmartOne (thesmartone):

You can not just get the answer. That is against teh Code of Conducts and can get someone suspended.

TheSmartOne (thesmartone):

@Student565

geerky42 (geerky42):

And yet many of us get away with that.

geerky42 (geerky42):

Plus I have seen mod giving only answer. Rules are not strict. @TheSmartOne

TheSmartOne (thesmartone):

@geerky42 but then again only the mods can suspend...

geerky42 (geerky42):

And because some of them don't ever follow rules, no need to worry about rules, really. Just do your thing... (unless you repeatedly break rule while it is clearly avoidable)

TheSmartOne (thesmartone):

thanks lol